OpenStudy (goformit100):

Find the equation of the straight line passing through (1, 2) and perpendicular to the line x+y+ 7 = 0

OpenStudy (dls):

m=1

OpenStudy (goformit100):

@dan815

OpenStudy (dls):

x=1 y=2 m=1

OpenStudy (dls):

y-2=1(x-1)

OpenStudy (goformit100):

How... please tell in details.... I can't understand

OpenStudy (dan815):

xD

OpenStudy (anonymous):

use y=mx+c form

OpenStudy (dan815):



OpenStudy (dan815):

say this is a given vector

OpenStudy (dan815):

imagine drawing a box around it of width a1 and height b1

OpenStudy (goformit100):

Ok

OpenStudy (dan815):